# Covariate-Adaptive Randomization in Clinical Trials Without Inflated Variances

### v1

@@ 2.2 Theoretical Properties
$\frac{\textsf{E}\big(\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\big)^{2}}{n}\to\vec{\sigma}_{Z}^{2}\text{ and }\frac{\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}}{\sqrt{n}}\overset{d}{\to}N(0,\vec{\sigma}_{Z}^{2}),$ (2.19)
where $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})])^{2}]$.
When $\gamma=0$, the CAR procedure is similar to Liu, Hu and Ma, (2025)’s procedure. It does not cause the “shift problem” when $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, and the variance is not inflated when $\rho=1/2$. When $\bm{X}$ is discrete as in Example 2.1 and $\bm{\phi}(\bm{X})$ given there with $w_{s}\neq 0$, i.e., the stratified randomization is adopted, the two conditions $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$ and $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)^{2}\in Span\{\bm{\phi}(\bm{X})\}$ are both satisfied for any random variables $Z$, and so the asymptotic normality (2.7) holds with a variance $\vec{\sigma}_{Z}^{2}$ being not inflated. This conclusion is consistent with Theorem 3 of Liu, Hu and Ma, (2025).
@@ 3 An application
    under $H_{A}:\tau\neq 0$, consider a sequence of local alternatives, i.e., $\tau=\delta/\sqrt{n}$ for a fixed $\delta\neq 0$ (we suppose the distributions $Y(t)-\mu_{t}$, $t=1,2$, do not depend on $n$), then
@@ 4.1 The properties of the CAR procedures
Let $\mathscr{F}_{n}=\sigma(\bm{X}_{n},\cdots,\bm{X}_{1},T_{n},\ldots,T_{1})$ be the history sigma generated by results of covariats and allocations up to stage, $\ell_{n}=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\Phi}_{n}\rangle}{(n-1)^{\gamma}}\right)$. Then, $\textsf{P}(T_{n+1}=1|\mathscr{F}_{n},\bm{X}_{n+1})=\ell_{n+1}$ by (2.1). Let $r\in[2,r_{0}]$. Write $\bm{\psi}_{n}=(T_{i}-\rho)\bm{\phi}(\bm{X}_{n})$. Then $\bm{\Lambda}_{n+1}=\bm{\Lambda}_{n}+\bm{\psi}_{n}$. From the elementary inequality that
@@
$\leq$ $-r\lambda_{1}\lambda_{2}\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)+c_{r}\left(\beta_{r}+\beta_{2}\|\bm{\Lambda}_{n}\|^{r-2}\right),$
@@
For (2.6), by noting that
$\frac{\textsf{E}\left|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\right|}{n}$
$\leq$ $\frac{\textsf{E}\left|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}I\{|Z_{i}|\leq M\}\right|}{n}+\frac{\sum_{i=1}^{n}|Z_{i}|I\{|Z_{i}|>M\}}{n}$
$\leq$ $\frac{\textsf{E}\left|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}I\{|Z_{i}|\leq M\}\right|}{n}+\textsf{E}[|Z|I\{|Z|>M\}],$ (4.34)
without loss of generality, we can assume that $\textsf{E}Z^{2}<\infty$.
Notice that $\Delta M_{i,Z}=:(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]$ is a sequence of martingale differences. It follows that
$\textsf{E}\left(\sum_{i=1}^{n}\left[(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]\right]\right)^{2}$
$=$ $\sum_{i=1}^{n}\textsf{E}\left(\left[(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]\right]\right)^{2}\leq n\textsf{E}[Z^{2}].$ (4.35)
On the other hand,
$\left|\textsf{E}\left[(\ell_{n+1}-\rho)Z_{n+1}\Big|\bm{\Lambda}_{n}]\right]\right|\leq C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|\|\bm{\Phi}_{n+1}\|}{n^{\gamma}}|Z_{n+1}|\Big|\bm{\Lambda}_{n}\right]$
$\leq$ $\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\big(\textsf{E}\|\bm{\Phi}\|^{2}\big)^{1/2}\big(\textsf{E}Z^{2}\big)^{1/2}\leq C\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}.$
Since, $\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}$, we have
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]=O\left(n^{\gamma/2}+n^{\frac{1}{4}+\frac{1}{4(r_{0}-1)}}\right)=o(n^{\gamma}).$
Thus,
$\textsf{E}\left[\left|\textsf{E}\left[(\ell_{n+1}-\rho)Z_{n+1}\Big|\bm{\Lambda}_{n}]\right]\right|\right]\leq C\frac{\textsf{E}[\|\bm{\Lambda}_{n}\|]}{n^{\gamma}}\to 0.$
It follows that
$\textsf{E}\left|\sum_{i=2}^{n}\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]\right|\leq C+C\sum_{i=2}^{n}\frac{\textsf{E}[\|\bm{\Lambda}_{i-1}\|]}{(i-1)^{\gamma}}=o(n).$
Hence, (2.6) is proven.
@@
and $\textsf{E}[W_{i}^{2}|\mathscr{F}_{i-1}]=\textsf{E}W^{2}$,
@@
$\textsf{E}\left[\Delta M_{i,\tilde{Z}}W_{i}\big|\mathscr{F}_{i-1}\right]=$ $\textsf{E}[(\ell_{i}-\rho)\tilde{Z}_{i}W_{i}|\bm{\Lambda}_{i-1}]\to 0\text{ in }L_{1}.$
@@
Suppose $\textsf{E}[Z^{2}]<\infty$. Then (2.2) follows from (2.8) and (4.38), and, (2.7) follows from (2.2) and (4.38). Notice
@@
By (4.38) and (2.8), (2.11) holds. At last, (2.6) follows from (4.1) and (2.7) when $\textsf{E}|Z|<\infty$. $\Box$
@@
Proof of Theorem 2.4. The proof of (i) is the same as that of (2.5) in Theorem 2.1. For (ii), suppose $\textsf{E}[Z|\bm{\phi}]=\langle\bm{z}_{0},\bm{\phi}\rangle\in Span\{\bm{\phi}\}$. Let $\widetilde{Z}=Z-\textsf{E}[Z|\bm{\phi}]$. Then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}+\langle\bm{z}_{0},\bm{\Lambda}_{n}\rangle,$
and $\{(T_{i}-\rho)\widetilde{Z}_{i}\}$ is a sequence of martingale differences with $|(T_{i}-\rho)\widetilde{Z}_{i}|\leq|Z_{i}|+\|\bm{y}_{0}\|\cdot\|\bm{\phi}_{i}\|$. Thus, the martingale differences are uniformly integrable which implies
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}}{n}\to 0\text{ in }L_{1}.$
(2.6) is proved by (i).
Suppose $\textsf{E}Z^{2}<\infty$. In this case $\langle\bm{z}_{0},\bm{\phi}\rangle=\textsf{Pj}[Z|\bm{\phi}]$ and thus $|\langle\bm{z}_{0},\bm{\Lambda}_{n}\rangle|\leq C_{0}(\textsf{E}Z^{2})^{1/2}\|\bm{\Lambda}_{n}\|$ as we have shown. For the martingale $\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}$, we have
@@
For (iii), notice that $\{(T_{i}-\rho)\widetilde{Z}_{i}\}$ is a sequence of martingale differences. By the central limit theorem(Hall and Heyde, , 1980, Theorem 3.2), it is sufficient to show that
@@
When $\rho=1/2$, then $(T_{i}-1/2)^{2}=1/4$, and (4.43) is obvious. When $\rho\neq 1/2$,
$(T_{i}-\rho)^{2}\widetilde{Z}_{i}^{2}=\rho(1-\rho)\widetilde{Z}_{i}^{2}+(1-2\rho)(T_{i}-\rho)\widetilde{Z}_{i}^{2}.$
Notice that $\textsf{E}[\widetilde{Z}^{2}|\bm{\phi}]\in Span\{\bm{\phi}\}$ by the assumption. Thus
$\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}^{2}\to 0\text{ in }L_{1}$
by (ii). Hence, (4.43) holds. The proof is completed. $\Box$

### v2

@@ 2.2 Theoretical Properties
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}}{\sqrt{n}}\overset{d}{\to}N(0,\vec{\sigma}_{Z}^{2}),$ (2.20)
where $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}]$. Further, if $\textsf{E}W^{2}<\infty$, $\textsf{E}[W]=0$ and also $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)W$ is in $Span\{\bm{\phi}(\bm{X})\}$, then we have (2.8).
When $\gamma=0$, the CAR procedure is similar to Liu, Hu and Ma, (2025)’s procedure. It does not cause the “shift problem” when $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, and the variance is not inflated when $\rho=1/2$. When $\bm{X}$ is discrete as in Example 2.1 and $\bm{\phi}(\bm{X})$ given there with $w_{s}\neq 0$, i.e., the stratified randomization is adopted, then $\textsf{E}[Z|\bm{\phi}(\bm{X})]$, $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)^{2}$, $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)W$ are all in $Span\{\bm{\phi}(\bm{X})\}$, and so the asymptotic normality (2.8) holds with a variance $\vec{\sigma}_{Z}^{2}$ being not inflated. This conclusion is consistent with Theorem 3 of Liu, Hu and Ma, (2025).
@@ 3 An application
    under $H_{A}:\tau\neq 0$, consider a sequence of local alternatives, i.e., $\tau=\delta/\sqrt{n}$ for a fixed $\delta\neq 0$ (we suppose the distributions of $Y(t)-\mu_{t}$, $t=1,2$, do not depend on $n$), then
@@
3.1 and 3.2 will follow from (2.7) and (2.8). When the stratified randomization is adopted, Theorems 3.1 and 3.2 remain true even when $\gamma=0$ by Theorem 2.4.
It is seen that the asymptotic variance $\vec{\sigma}_{Z}^{2}$ in (2.8) has the closed form $\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$ and does not exceed $\textsf{E}Z^{2}$ is important because in this case the classical test is valid and the variance is estimable so that the test can be adjusted to have the precise type I error rate.
When the stratified randomization is adopted, we always have $\vec{\sigma}_{Z}^{2}=\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$. This is the main reason why the classical test under the stratified randomization is always valid and can be adjusted in varies ways.
@@ 4.1 The properties of the CAR procedures
Let $\mathscr{F}_{n}=\sigma(\bm{X}_{i},T_{i},Z_{i},W_{i},i=1,2,\ldots,n)$ be the history sigma generated by results of variables and allocations up to stage, $\ell_{n}=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\Phi}_{n}\rangle}{(n-1)^{\gamma}}\right)$. Then, $\textsf{P}(T_{n+1}=1|\mathscr{F}_{n},\bm{X}_{n+1})=\ell_{n+1}$ by (2.1). Let $r\in[2,r_{0}]$. Write $\bm{\psi}_{n}=(T_{i}-\rho)\bm{\phi}(\bm{X}_{n})$. Then $\bm{\Lambda}_{n+1}=\bm{\Lambda}_{n}+\bm{\psi}_{n}$. From the elementary inequality that
@@
$\leq$ $-r\lambda_{1}\lambda_{2}\|\bm{\Lambda}_{n}\|^{r-1}\left(\frac{\|\bm{\Lambda}_{n}\|}{n^{\gamma}}\wedge 1\right)+c_{r}\left(\beta_{r}+\beta_{2}\|\bm{\Lambda}_{n}\|^{r-2}\right),$ (4.31)
@@
Notice that
$\textsf{E}|(\ell_{n}-1)Z_{n}|\leq$ $cM\textsf{E}\left[\frac{\|\bm{\Lambda}_{n-1}\|\cdot\|\bm{\phi}_{n}\|}{(n-1)^{\gamma}}\right]+\textsf{E}[|Z|I\{|Z|>M\}]$
$=$ $cM\frac{\textsf{E}\left[\|\bm{\Lambda}_{n-1}\|\right]\textsf{E}[\|\bm{\phi}\|]}{(n-1)^{\gamma}}+\textsf{E}[|Z|I\{|Z|>M\}]$
$\to$ $0\text{ as }n\to\infty\text{ and then }M\to\infty,$
since
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|\right]=O\left(n^{\gamma/2}+n^{\frac{1}{4}+\frac{1}{4(r_{0}-1)}}\right)=o(n^{\gamma}),\;\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}.$
(2.6) is proven.
For (2.7), let $\Delta M_{i,Z}=:(T_{i}-\rho)Z_{i}-\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]$, $M_{n,Z}=\sum_{i=1}^{n}\Delta M_{i,Z}$. Notice that $\{\Delta M_{i,Z},\mathscr{F}_{i},i=1,2,\ldots\}$ is a sequence of martingale differences. Since $|\Delta M_{i,Z}|\leq|Z_{i}|+\textsf{E}|Z|$, $\{\Delta M_{i,Z},i=1,2,\ldots\}$ is uniformly integrable. Thus, $\frac{1}{n}\textsf{E}[|M_{n,Z}|\to 0$ (Hall and Heyde, , 1980, Theorem 2.2). On the other hand,
$\textsf{E}\Big|\sum_{i=2}^{n}\textsf{E}[(\ell_{i}-\rho)Z_{i}|\bm{\Lambda}_{i-1}]\Big|\leq\sum_{i=2}^{n}\textsf{E}[|(\ell_{i}-\rho)Z_{i}|]=o(n)$
by (2.6). Hence, (2.7) is proven.
@@
i.e., the Linderberg condition is satisfied. Further, and $\textsf{E}[W_{i}^{2}|\mathscr{F}_{i-1}]=\textsf{E}W^{2}$,
@@
$\textsf{E}\left[\Delta M_{i,\tilde{Z}}W_{i}\big|\mathscr{F}_{i-1}\right]=$ $\textsf{E}[(\ell_{i}-\rho)\tilde{Z}_{i}W_{i}|\bm{\Lambda}_{i-1}]\to 0\text{ in }L_{1},$
by applying (2.6) to $\tilde{Z}^{2}_{n}$ and $\tilde{Z}_{n}W_{n}$.
@@
When $\textsf{E}|Z|<\infty$, the proofs of (2.6) and (2.7) are the same as those in Theorem 2.1 since $\textsf{E}\|\bm{\Lambda}_{n}\|=o(n^{\gamma})$.
Suppose $\textsf{E}[Z^{2}]<\infty$. Then ((ii)) follows from (2.9) and (4.38), and, (2.8) follows from ((ii)) and (4.38). Notice
@@
By (4.38) and (2.9), (2.12) holds. $\Box$
@@
Proof of Theorem 2.4. For (i), notice that (4.1) still holds for $\gamma=0$. When $\|\bm{\Lambda}_{n}\|\geq\frac{2c_{r}\beta_{2}}{r\lambda_{1}\lambda_{2}}+1$,
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}+c_{r}\beta_{r}.$
When $\|\bm{\Lambda}_{n}\|\leq\frac{2c_{r}\beta_{2}}{r\lambda_{1}\lambda_{2}}+1$, $\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}+C_{r}$. Thus, exist $a_{r},b_{r}>0$ such that $\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-a_{r}\|\bm{\Lambda}_{n}\|^{r-1}+b_{r}.$ It follows that
$\textsf{E}[\|\bm{\Lambda}_{n+1}\|^{r}]-\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq$ $-a_{r}\textsf{E}[\|\bm{\Lambda}_{n}\|^{r-1}]+b_{r}$ (4.43)
$\leq$ $b_{r}.$ (4.44)
From (4.44), it follows that $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r_{0}}]\leq nb_{r_{0}}$.
Now, let $m:=m_{n}\in\{0,1,\ldots,n\}$ be the last one for which $a_{r}\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1}]+b_{r}\geq 0.$ Let $r-1+q=r_{0}$, $1/p+1/q=1$. Then
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq$ $\textsf{E}\left[\|\bm{\Lambda}_{m+1}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r}\right]+b_{r}=b_{r}+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{\frac{r-1}{p}}\|\bm{\Lambda}_{m}\|^{\frac{r-1}{q}+1}\right]$
$\leq$ $b_{r}+\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1}\right]\right)^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1+q}\right]\right)^{1/q}$
$\leq$ $b_{r}+(b_{r}/a_{r})^{1/p}\left(b_{r_{0}}m\right)^{1/q}\leq C_{r,r_{0}}n^{1/q}.$ (4.45)
Letting $r=2$ and $q=r_{0}-1$ yields $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{\frac{1}{r_{0}-1}})$. When $r_{0}\geq 3$. Let $r=(1+r_{0})/2$. Then $r\geq 2$, $q=r$. Thus, $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]=O(n^{1/r})$, and so $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{2/r^{2}})=O(n^{\frac{8}{(r_{0}+1)^{2}}}$. The proof of (i) is completed.
For (ii), suppose $\textsf{E}[Z|\bm{\phi}]=\langle\bm{\phi},\bm{z}_{0}\rangle\in Span\{\bm{\phi}\}$. Let $\widetilde{Z}=Z-\textsf{E}[Z|\bm{\phi}]$. Then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}+\langle\bm{\Lambda}_{n},\bm{z}_{0}\rangle,$
and $\{(T_{i}-\rho)\widetilde{Z}_{i},\mathscr{F}_{i},i=1,2,\ldots\}$ is a sequence of martingale differences with $|(T_{i}-\rho)\widetilde{Z}_{i}|\leq|Z_{i}|+\|\bm{z}_{0}\|\cdot\|\bm{\phi}_{i}\|$. Thus, the martingale differences are uniformly integrable which implies
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}}{n}\to 0\text{ in }L_{1}$
(Hall and Heyde, , 1980, Theorem 2.22). (2.7) is proved by (i).
Suppose $\textsf{E}Z^{2}<\infty$. In this case $\langle\bm{\phi},\bm{z}_{0}\rangle=\textsf{Pj}[Z|\bm{\phi}]$ and thus $|\langle\bm{\Lambda}_{n},\bm{z}_{0}\rangle|\leq C_{0}(\textsf{E}Z^{2})^{1/2}\|\bm{\Lambda}_{n}\|$ as we have shown. For the martingale $\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}$, we have
@@
For (iii), by the central limit theorem(Hall and Heyde, , 1980, Theorem 3.2), it is sufficient to show that
@@
Notice
$(T_{i}-\rho)^{2}\widetilde{Z}_{i}^{2}=\rho(1-\rho)\textsf{E}\widetilde{Z}^{2}+\rho(1-\rho)(\widetilde{Z}_{i}^{2}-\textsf{E}\widetilde{Z}^{2})+(1-2\rho)(T_{i}-\rho)\widetilde{Z}_{i}^{2}.$
It is obvious that
$\frac{\sum_{i=1}^{n}(\widetilde{Z}_{i}^{2}-\textsf{E}\widetilde{Z}^{2})}{n}\to 0\text{ in }L_{1}.$
When $\rho\neq 1/2$, we have
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}^{2}}{n}\to 0\text{ in }L_{1},$
by (ii) and noticing the assumption that $\textsf{E}[\widetilde{Z}^{2}|\bm{\phi}]\in Span\{\bm{\phi}\}$ . Hence, (4.46) holds.
When $(Z-\textsf{E}[Z|\bm{\phi}])W$ is also in $Span\{\bm{\phi}\}$. Then
$\frac{\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{Z}_{i}W_{i}}{n}\to 0\text{ in }L_{1}$
by (ii) again. Thus,
$\sum_{i=1}^{n}\big((T_{i}-\rho)\widetilde{Z}_{i},W_{i}\big)^{\otimes 2}\to diag\big(\vec{\sigma}_{Z}^{2},\textsf{E}W^{2}\big)\text{ in }L_{1}.$
Applying the central limit theorem to martingale differences $\Big\{\big((T_{i}-\rho)\widetilde{Z}_{i},W_{i}\big),\mathscr{F}_{i},i=1,2,\ldots\Big\}$ yields (2.8). The proof is completed. $\Box$